# QNPU: Quantum Network Processor Unit for Quantum Supercomputers

### v2

@@ 1. Introduction
As quantum computing advances, its potential to solve large-scale problems in domains such as cryptography (Shor, 1994), and chemistry problems (McArdle et al., 2020) becomes increasingly apparent. These large-scale problems often require computational resources far beyond the capacity of a single quantum processor. One promising solution for large-scale problems is distributed quantum computing (DQC) (Cuomo et al., 2020; Caleffi et al., 2024), which enables scalability by partitioning large quantum algorithms across multiple quantum processors, each contributing a portion of the overall computation.
@@
Significant efforts have been devoted toward realizing DQC. Some works (Cuomo et al., 2020; Rodrigo et al., 2020; Ang et al., 2024) propose full-stack layered architectures to capture the broad challenges of building distributed quantum systems. Others target specific challenges: At the software level, frameworks such as the Quantum Message Passing Interface (QMPI) (Häner et al., 2021) extend the classical MPI standard to quantum systems, providing programming models for DQC. From the compiler level, a variety of frameworks (Wu et al., 2022b, a, 2023) have been developed to optimize the scheduling of inter-node communication, thereby reducing the overhead of distributed execution. On the hardware side, demonstrations across different qubit technologies, including superconducting qubits (Magnard et al., 2020; Smith et al., 2022; Storz et al., 2023), ion traps (Monroe et al., 2014) have showcased the feasibility of multi-core and multi-node platforms.
@@
However, despite significant progress in demonstrating DQC capabilities in multi-core and multi-node platforms, current systems lack specialized hardware to efficiently process inter-node communications. When distributed quantum algorithms require frequent communication between processors, existing approaches suffer from substantial performance bottlenecks due to the absence of dedicated communication processing units. This limitation becomes particularly severe for communication-intensive quantum algorithms, where the overhead of managing inter-node communications can dominate overall execution time.
To address this critical gap, we introduce the concept of a quantum supercomputer, a computing system that integrates multiple quantum processors through quantum and classical links to enable large-scale quantum computation. As a key component of this architecture, we propose the Quantum Network Processor Unit (QNPU), a specialized hardware unit designed to efficiently execute quantum communication protocols between processors.
@@
As shown in Figure 1, a quantum supercomputer is organized as interconnected quantum nodes. Each node has a Quantum Processor Unit (QPU) that executes local quantum computations within the quantum node, and a Quantum Network Processor Unit (QNPU) that performs quantum communication protocols between quantum nodes. Both the QPU and QNPU consist of two key components: a classical controller managing instruction execution; and a quantum chip containing physical qubits for quantum operations. Within each quantum node, QPU and QNPU interact through both a classical link and a quantum link. Intra-node classical links connect their controllers, enabling the QPU to delegate quantum communication protocols to the QNPU for execution; Intra-node quantum links connect their quantum chips, enabling qubit movement between the QPU and the QNPU. Communication between quantum nodes involves inter-node classical links and quantum links. Inter-node classical links support classical data transfer between nodes; inter-node quantum links support building the entanglement between nodes.
@@
This decoupled architecture, which incorporates QPU and QNPU in one quantum node and separates local computation (handled by QPUs) from quantum communication protocols (managed by QNPUs), enables efficient execution of distributed quantum algorithms.
@@
Proposed a decoupled architecture for quantum supercomputers, which separates local computation (handled by QPUs) from quantum communication protocols (managed by QNPUs).
@@
By bridging the gap between distributed quantum algorithms and quantum supercomputer hardware, our work paves the way for scalable and efficient quantum computing, bringing us closer to realizing the full potential of quantum computing.
@@ 2.1. EPR Pair Generation
Inter-node Quantum communication in a quantum supercomputer consumes entangled states between nodes for quantum state and gate teleportation (Wan et al., 2019; Chou et al., 2018; Hermans et al., 2022). One of the widely used resource entangled states is Einstein–Podolsky
–Rosen (EPR) pairs (Einstein et al., 1935), which is represented as .
In contrast to quantum network applications, where EPR pairs are generated between remote nodes separated by long distances, typically on the order of kilometers (Wei et al., 2022; Knaut et al., 2024; Stolk et al., 2024), we focus on distributed quantum algorithms running on the quantum supercomputers, where communication nodes are located within a relatively short range, such as within a single room. This proximity makes the generation of EPR pairs significantly less challenging and time-consuming compared to quantum networks.
For instance, in superconducting quantum computing systems, the nodes of a quantum supercomputer consist of superconducting chips (Smith et al., 2022; Guinn et al., 2023). The physical connections enabling EPR pair generation can be coaxial cables that directly link superconducting qubits across different nodes—an approach that has already been experimentally demonstrated (Magnard et al., 2020; Storz et al., 2023). By cooling these coaxial cables to cryogenic temperatures, the loss of microwave photon propagation can be minimized. Consequently, EPR pairs between nodes can be efficiently generated through the direct exchange of microwave photons (Storz et al., 2023; Campagne-Ibarcq et al., 2018; Besse et al., 2020; Niu et al., 2023; Grebel et al., 2024).
@@ 2.2. EPR Pair Buffering
Efficient utilization of communication resources in quantum nodes requires a strategy for handling EPR pairs after generation. The concept of EPR pair buffering, proposed in (Wu et al., 2022a), addresses this need by introducing a buffering mechanism to store generated EPR pairs. As illustrated in Fig. 2, once an EPR pair is generated on the communication qubits of two connected quantum nodes, the entangled state is swapped into other available data qubits in the quantum node. This process releases the communication qubits, enabling them to generate new EPR pairs. The buffered EPR pairs remain available for future quantum communication tasks.
@@ 2.3. Quantum Communication Protocols
With the shared EPR pairs between quantum nodes, two types of quantum communication protocols, TP-Comm and Cat-Comm, can be utilized for quantum communication and remote gate operations (Häner et al., 2021; Wu et al., 2022a, b, 2023).
Figure 3 illustrates example quantum circuits for these protocols. The TP-Comm protocol, shown in Fig. 3(a), employs a prepared EPR pair to teleport a qubit state from node A to node B. Once the state is teleported, the qubit in node B can serve as the control qubit in a CNOT gate, effectively realizing a remote CNOT gate between the two nodes.
In comparison, the Cat-Comm protocol, shown in Fig.3(b), utilizes a shared EPR pair to entangle the qubit in node B with the data qubit in node A through the “cat-entangler” circuit. After generating this entanglement, the qubit in node B acts as an intermediary for the remote CNOT operation. Subsequently, the “cat-disentangler” circuit disentangles the qubit in node B from the other two qubits to complete the remote CNOT operation. Unlike the TP-Comm protocol, the Cat-Comm protocol does not teleport the quantum state, which restricts the supported remote gate operations. However, Cat-Comm enables “broadcasting” of entanglement, making it suitable for multi-node or collective operations(Häner et al., 2021; Wu et al., 2022a).
@@ 2.4. Related Work
Quantum networks: Quantum networks can be classified into two types. The first type, quantum internet (Pirandola and Braunstein, 2016; Wehner et al., 2018), enables long-range quantum communication over optical links. These networks rely on quantum repeaters to mitigate photon loss. The second type, quantum intranet (Chow, 2021), enables quantum information transfer within shorter-range networks. Experimental work (Magnard et al., 2020; Storz et al., 2023) has demonstrated that superconducting qubits connected via microwave links can realize such short-range quantum networks. Our work focus on quantum supercomputers where each node in the quantum supercomputer is connected via short range microwave links.
@@
DQC Architectures: Several scalable architectural directions have been explored to realize DQC. The chiplet architecture (Smith et al., 2023; LaRacuente et al., 2025) integrates multiple smaller quantum chips via microwave links, analogous to multi-core classical systems where multiple processing cores collaborate to solve a problem in a distributed fashion. Another scalable design is the quantum data center (QDC) (Shapourian et al., 2025), which utilizes optical switches to interconnect multiple quantum processors. QDCs operate over short distances, therefore avoid the need for quantum repeaters, making them well suited with near-term hardware capabilities (Zhang et al., 2025). The distinction between these approaches lies in their scaling philosophy: chiplet architectures achieve performance improvements by integrating multiple quantum processing elements within a single node (scale-up), while QDC architectures enhance computational capacity by interconnecting multiple independent quantum nodes (scale-out). Our work follows the scale-out paradigm, focusing specifically on accelerating quantum communication between nodes through the design of specialized hardware, the Quantum Network Processor Unit (QNPU).
@@
DQC Compilers: A variety of compiler frameworks (Wu et al., 2022b, a, 2023) have been proposed to optimize inter-node communications in distributed quantum systems. Other works focus on specific architectrue setting: Zhang et al.(Zhang et al., 2025) presents a compiler framework for EPR scheduling to mitigate entanglement generation latency in quantum data centers, and Zhang et al. (Zhang et al., 2024) presents Multi-Entry Communication Highway (MECH), which leverages compiler-managed ancillary qubits to enhance concurrency in chiplet architectures. While these efforts improve communication efficiency through software-level optimizations, our work introduces hardware acceleration via the Quantum Network Processing Unit (QNPU), a design that both complements and benefits from compiler-driven strategies.
@@
Instruction Set Architecture and Microarchitecture for Quantum Communications: QuMIS (Fu et al., 2017) and eQASM (Fu et al., 2019) are executable ISAs designed for quantum processors, focusing on local quantum operations but lacking support for quantum communication between distributed processors. NetQASM (Dahlberg et al., 2022) extends this by supporting quantum network applications with instructions for generating entanglement between remote quantum nodes. More recently, Donne et al. (Donne et al., 2025) introduced an architecture and operating system (QNodeOS) to support quantum network applications with multitasking capabilities. Our work builds upon these foundations but takes a fundamentally different approach by physically decoupling computation and communication processing units. By introducing a dedicated Quantum Network Processing Unit (QNPU) that operates in parallel with the Quantum Processing Unit (QPU), the inter-node communication can be executed more effeciently.
@@
Large-Scale Fault-Tolerant Distributed Quantum computer: Kim et al. (Kim et al., 2024) address the challenges of fault tolerance in million-qubit-scale quantum computers distributed across multiple dilution refrigerators. Their work focuses on implementing fault-tolerant quantum computation using surface codes with lattice surgery, with an emphasis on minimizing logical error rates in large-scale distributed systems. In contrast, our approach tackles the complementary challenge of providing efficient quantum communication between quantum nodes through a dedicated Quantum Network Processing Unit (QNPU). Kim et al.’s fault-tolerant techniques could be integrated with our QNPU-based communication stack to construct a scalable distributed quantum system that is both fault-tolerant and communication-efficient.
@@ 3.1. Layer Stack for Quantum Supercomputers
Inspired by the classical network architecture’s Open Systems Interconnection (OSI) model, previous works have proposed layered quantum network stacks to support quantum network applications (Dahlberg et al., 2019; Kozlowski et al., 2020). Similarly, the work in (Ang et al., 2024) introduced a layer stack for a multinode quantum computer (MNQC) architecture. Our proposed layer stack for quantum supercomputers draws upon these ideas while tailoring the design for the efficient execution of distributed quantum algorithms. Figure 4 illustrates our four-layer architecture.
Physical Layer: The physical layer forms the foundation of the quantum supercomputer. It is responsible for generating raw EPR pairs between directly connected quantum nodes. As our design adopts the established protocols described in (Dahlberg et al., 2019; Kozlowski et al., 2020) for the physical layer, we do not further elaborate on its specifics here.
Data Link Layer: We define the data link layer’s primary function as EPR prefetch, which ensures reliable pre-generation and buffering (Wu et al., 2022a) of required EPR resources to hide EPR pair generation latency. Prior work (Zhang et al., 2025) has proposed compiler frameworks enables look-ahead scheduling of EPR pairs, which can be integrated into the data link layer to realize this prefetch functionality.
@@ 3.2. Decoupled Processing Units for Quantum Supercomputers
As established in Section 3.1, our architecture employs a clear separation of concerns: the application layer hosts Quantum Processing Unit (QPU) for executing distributed quantum algorithms, when the QPU encounters operations requiring quantum communication, it offloads these tasks to the QNPU. This decoupled design contrasts with traditional designs (Dahlberg et al., 2022) that consolidate both computation and communication functions within a single processing unit.
In our design as shown in Figure 5, the QPU’s qubits constitute the ”computation zone,” dedicated solely to executing local quantum operations, whereas the QNPU’s qubits form the ”communication zone,” optimized for performing quantum communication protocols. This zone-based organization requires transitions between zones for quantum states that need to participate in both computation and communication operations. Works by Bluvstein et al. (Bluvstein et al., 2022, 2023) demonstrated the feasibility of such zone architectures and the required qubit movement capabilities. Their implementations of zone-based qubit manipulation enable efficient state transfer between distinct functional regions, providing strong experimental validation for our decoupled architectural approach. These advances in quantum hardware directly support the physical realizability of our design decisions.
The advantages of this decoupled architecture are multifold. First, it allows specialized optimization of each processing unit for its specific task, improving both computational and communication efficiency. Second, it enables parallel operation of computation and communication tasks, reducing overall execution time for distributed quantum algorithms. Third, it provides a clean separation of concerns that simplifies system design, implementation, and future scaling. These benefits directly address the performance bottlenecks in distributed quantum computing, particularly for communication-intensive algorithms like those evaluated in Section 7.

### v3

# A Quantum Network Processor Unit for Distributed Quantum Computing

@@ 1. Introduction
As quantum computing advances, its potential to solve large-scale problems in domains such as cryptography (shor1994factor), and chemistry problems (McArdle2020chemistry) becomes increasingly apparent. These large-scale problems often require computational resources far beyond the capacity of a single quantum processor. One promising solution for large-scale problems is distributed quantum computing (DQC) (Cuomo2020DQC; Caleffi2024DQC), which enables scalability by partitioning large quantum algorithms across multiple quantum processors, each contributing a portion of the overall computation.
@@
Significant efforts have been devoted toward realizing DQC. Some works (Cuomo2020DQC; Rodrigo2020DQC; ang2024ARQUIN) propose full-stack layered architectures to capture the broad challenges of building distributed quantum systems. Others target specific challenges: At the software level, frameworks such as the Quantum Message Passing Interface (QMPI) (haner2021QMPI) extend the classical MPI standard to quantum systems, providing programming models for DQC. From the compiler level, a variety of frameworks (wu2022autocomm; wu2022collcomm; Wu2023QuComm) have been developed to optimize the scheduling of inter-node communication, thereby reducing the overhead of distributed execution. On the hardware side, demonstrations on various qubit technologies, including superconducting qubits (Magnard2020microwave; Smith2022Chiplet; Storz2023), ion traps (Main2025dqc) have showcased the feasibility of distributed quantum systems.
Looking forward, the quantum computing community increasingly recognizes that the future of quantum computing lies not in a single technology, but in the strategic integration of diverse qubit types since no single qubit technology excels at all computational tasks. Recent initiatives such as the Defense Advanced Research Projects Agency (DARPA)’s Heterogeneous Architectures for Quantum (HARQ) program (DARPA_HARQ_2025) seeks to explore heterogeneous quantum systems to realize DQC. In addition, heterogeneous quantum architectures require specialized infrastructure capable of managing communication across different species of qubits.
@@
Even within current homogeneous distributed quantum systems, existing architectures lack the specialized hardware needed for efficient inter-node communication. When distributed quantum algorithms require frequent communication between processors, existing approaches suffer from substantial performance bottlenecks due to the absence of dedicated communication processing units. This limitation becomes particularly severe for communication-intensive quantum algorithms, where the overhead of managing inter-node communications can dominate overall execution time. For future heterogeneous systems, this challenge is further compounded by the need to transfer quantum states between qubits with different species.
To support both homogeneous and heterogeneous distributed quantum computing systems, we introduce the concept of a quantum supercomputer: a scalable computing system composed of multiple quantum processors interconnected through quantum and classical links. At the core of this architecture is the Quantum Network Processor Unit (QNPU), a specialized hardware module designed to efficiently execute quantum communication protocols. By offloading communication tasks from the Quantum Processing Unit (QPU), the QNPU provides a clean separation between local computation and quantum communication, enabling efficient state transfer and coordination across processors. While our evaluation focuses on homogeneous systems, the same decoupled design naturally extends to heterogeneous quantum platforms. In such systems, different qubit technologies or hardware zones can interoperate through the QNPU, making it a foundational building block for future heterogeneous quantum computers.
@@
As shown in Fig. 1, a quantum supercomputer is organized as interconnected quantum nodes. Each node has a Quantum Processor Unit (QPU) that executes local quantum computations within the quantum node, and a Quantum Network Processor Unit (QNPU) that performs quantum communication protocols between quantum nodes. Both the QPU and QNPU consist of two key components: a classical controller managing instruction execution; and a quantum device containing physical qubits for quantum operations. Within each quantum node, QPU and QNPU interact through both a classical link and a quantum link. Intra-node classical links connect their controllers, enabling the QPU to delegate quantum communication protocols to the QNPU for execution; Intra-node quantum links connect their quantum devices, enabling qubit movement between the QPU and the QNPU. Communication between quantum nodes involves inter-node classical links and quantum links. Inter-node classical links support classical data transfer between nodes; inter-node quantum links support building the entanglement between nodes.
@@
This decoupled architecture, which incorporates QPU and QNPU in one quantum node and separates local computation (handled by QPUs) from quantum communication protocols (managed by QNPUs), enables efficient execution of distributed quantum algorithms. By abstracting communication protocols away from the QPU implementation, this architecture also provides flexibility for future heterogeneous configurations where different nodes may employ different qubit technologies.
@@
Proposed a decoupled architecture for quantum supercomputers, which separates local computation (handled by QPUs) from quantum communication protocols (managed by QNPUs), providing a foundation for both homogeneous and heterogeneous distributed quantum architectures.
@@ 2.1. EPR Pair Generation
Inter-node quantum communication in a quantum supercomputer consumes entangled states between nodes for quantum state and gate teleportation (Wan2019; Chou2018; Hermans2022). One of the widely used resource entangled states is Einstein–Podolsky–Rosen (EPR) pairs (Einstein1935epr), represented as .
In contrast to quantum network applications, where EPR pairs are generated between remote nodes separated by long distances, typically on the order of kilometers (Wei2022; Knaut2024; Stolk2024), we focus on distributed quantum algorithms running on the quantum supercomputers, where communication nodes are located within a relatively short range, such as within a single room. This proximity makes the generation of EPR pairs significantly less challenging and time-consuming compared to quantum networks.
For instance, in superconducting quantum computing systems, the nodes of a quantum supercomputer consist of superconducting chips (Smith2022Chiplet; guinn2023Quirc). The physical connections enabling EPR pair generation can be coaxial cables that directly link superconducting qubits across different nodes—an approach that has already been experimentally demonstrated (Magnard2020microwave; Storz2023). By cooling these coaxial cables to cryogenic temperatures, the loss of microwave photon propagation can be minimized. Consequently, EPR pairs between nodes can be efficiently generated through the direct exchange of microwave photons (Storz2023; Campagne-Ibarcq2018; Besse2020; Niu2023interconnect; Grebel2024).
@@ 2.2. EPR Pair Buffering
Efficient utilization of communication resources in quantum nodes requires a strategy for handling EPR pairs after generation. The concept of EPR pair buffering, proposed in (wu2022collcomm), addresses this need by introducing a buffering mechanism to store generated EPR pairs. As illustrated in Fig. 2, once an EPR pair is generated on the communication qubits (marked with orange color) of two connected quantum nodes, the entangled state is swapped into other available data qubits (marked with blue color) in the quantum node. This process releases communication qubits, enabling them to generate new EPR pairs. The buffered EPR pairs remain available for future quantum communication tasks.
@@ 2.3. Quantum Communication Protocols
With shared EPR pairs between quantum nodes, two quantum communication protocols, TP-Comm and Cat-Comm, can be utilized for quantum communication and remote gate operations (haner2021QMPI; wu2022collcomm; wu2022autocomm; Wu2023QuComm).
Fig. 3 shows the quantum circuits for these protocols. The TP-Comm protocol, shown in Fig. 3(a), employs a prepared EPR pair to teleport a qubit state from node A to node B. Once the state is teleported, the qubit in node B can serve as the control qubit in a CNOT gate, effectively realizing a remote CNOT gate between the two nodes.
In comparison, the Cat-Comm protocol, shown in Fig.3(b), utilizes a shared EPR pair to entangle the qubit in node B with the data qubit in node A through the “cat-entangler” circuit. After generating this entanglement, the qubit in node B acts as an intermediary for the remote CNOT operation. Subsequently, the “cat-disentangler” circuit disentangles the qubit in node B from the other two qubits to complete the remote CNOT operation. Unlike the TP-Comm protocol, the Cat-Comm protocol does not teleport the quantum state, which restricts the supported remote gate operations. However, Cat-Comm enables “broadcasting” of entanglement, making it suitable for multi-node or collective operations(haner2021QMPI; wu2022collcomm).
@@ 2.4. Related Work
Quantum Networks: Quantum networks can be classified into two types. The first type, quantum internet (Pirandola2016internet; Stephanie2018internet), enables long-range quantum communication over optical links. These networks rely on quantum repeaters to mitigate photon loss. The second type, quantum intranet (chow2021DQC), enables quantum information transfer within shorter-range networks. Experimental work (Magnard2020microwave; Storz2023) has demonstrated that superconducting qubits connected via microwave links can realize such short-range quantum networks. Our work focus on quantum supercomputers where each node in the quantum supercomputer is connected via short range microwave links.
@@
DQC Architectures: Several scalable architectural directions have been explored to realize DQC. The chiplet architecture (Smith2022Chiplet; LaRacuente_2025) integrates multiple smaller quantum chips via microwave links, analogous to multi-core classical systems where multiple processing cores collaborate to solve a problem in a distributed fashion. Another scalable design is the quantum data center (QDC) (shapourian2025quantumdatacenterinfrastructures), which utilizes optical switches to interconnect multiple quantum processors. QDCs operate over short distances, therefore avoid the need for quantum repeaters, making them well suited with near-term hardware capabilities (zhang2024QDC). The distinction between these approaches lies in their scaling philosophy: chiplet architectures achieve performance improvements by integrating multiple quantum processing elements within a single node (scale-up), while QDC architectures enhance computational capacity by interconnecting multiple independent quantum nodes (scale-out). Our work follows the scale-out paradigm, focusing specifically on accelerating quantum communication between nodes through the design of specialized hardware, the Quantum Network Processor Unit (QNPU).
@@
DQC Compilers: A variety of compiler frameworks (wu2022autocomm; wu2022collcomm; Wu2023QuComm) have been proposed to optimize inter-node communications in distributed quantum systems. Other works focus on specific architecture setting: Zhang et al.(zhang2024QDC) presents a compiler framework for EPR scheduling to mitigate entanglement generation latency in quantum data centers, and Zhang et al. (Zhang_2024_MECH) presents Multi-Entry Communication Highway (MECH), which leverages compiler-managed ancillary qubits to enhance concurrency in chiplet architectures. While these efforts improve communication efficiency through software-level optimizations, our work introduces hardware acceleration via the Quantum Network Processing Unit (QNPU), a design that both complements and benefits from compiler-driven strategies.
@@
Instruction Set Architecture and Microarchitecture for Quantum Communications: QuMIS (Fu2017QuMA) and eQASM (Fu2019eQASM) are executable ISAs designed for quantum processors, focusing on local quantum operations but lacking support for quantum communication between distributed processors. NetQASM (Dahlberg2022NetQASM) extends this by supporting quantum network applications with instructions for generating entanglement between remote quantum nodes. More recently, Donne et al. (Donne2025OS) introduced an architecture and operating system (QNodeOS) to support quantum network applications with multitasking capabilities. Both the NetQASM and QNodeOS introduce the concept of Quantum Network Processing Unit (QNPU) for executing applications on quantum network nodes, and their QNPUs are a unified processing unit that processes both local quantum operations and remote quantum operations. In contrast, our work takes a fundamentally different approach by physically decoupling computation and communication processing units, where QPUs are responsible for local computation and QNPUs are responsible for quantum communication operations.
@@
Large-Scale Fault-Tolerant Distributed Quantum Computer: Kim et al. (Kim2024MQDQC) address the challenges of fault tolerance in million-qubit-scale quantum computers distributed across multiple dilution refrigerators. Their work focuses on implementing fault-tolerant quantum computation using surface codes with lattice surgery, with an emphasis on minimizing logical error rates in large-scale distributed systems. In comparison, our approach tackles the complementary challenge of providing efficient quantum communication between quantum nodes through a dedicated Quantum Network Processing Unit (QNPU). Their proposed fault-tolerant techniques could be integrated with our QNPU-based communication stack to construct a scalable distributed quantum system that is both fault-tolerant and communication-efficient.
@@ 3.1. Layer Stack for Quantum Supercomputers
Inspired by the classical network architecture’s Open Systems Interconnection (OSI) model, previous works have proposed layered quantum network stacks to support quantum network applications (Dahlberg2019linklayer_protocol; Kozlowski_2020). Similarly, the work in (ang2024ARQUIN) introduced a layer stack for a multinode quantum computer (MNQC) architecture. Our proposed layer stack for quantum supercomputers draws upon these ideas while introduces dedicated hardware acceleration for communication protocols (through the QNPU at the network layer), proactive EPR resource management (EPR prefetch at the data link layer), and a clear architectural separation between local quantum computation and inter-node communication to enable efficient execution of distributed quantum algorithms. Fig. 4 illustrates our four-layer architecture.
Physical Layer: The physical layer forms the foundation of the quantum supercomputer. It is responsible for generating raw EPR pairs between directly connected quantum nodes. As our design adopts the established protocols (Dahlberg2019linklayer_protocol; Kozlowski_2020) for the physical layer, we do not further elaborate on its specifics.
Data Link Layer: We define the data link layer’s primary function as EPR prefetch, which ensures reliable pre-generation and buffering (wu2022collcomm) of required EPR resources to hide EPR pair generation latency. Prior work (zhang2024QDC) has proposed compiler frameworks enables look-ahead scheduling of EPR pairs, which can be integrated into the data link layer to realize this prefetch functionality.
@@ 3.2. Decoupled Processing Units for Quantum Supercomputers
As established in Section 3.1, our architecture employs a clear separation of concerns: the application layer hosts Quantum Processing Unit (QPU) for executing distributed quantum algorithms, when the QPU encounters operations requiring quantum communication, it offloads these tasks to the QNPU. This decoupled design contrasts with traditional designs (Dahlberg2022NetQASM) that consolidate both computation and communication functions within a single processing unit.
In our design as shown in Fig. 5, the QPU’s qubits constitute the ”computation zone,” dedicated solely to executing local quantum operations, whereas the QNPU’s qubits form the ”communication zone,” optimized for performing quantum communication protocols. This decoupled architecture offers significant advantages for both homogeneous and heterogeneous quantum computing scenarios. By separating computation from communication, QPU and QNPU can be specialized and optimized for their specific tasks. The superscalar QNPU pipeline design proposed in Section VI does not affect the QPU, as a result of this decoupled architecture. Moreover, by decoupling the QPU from network communication responsibilities, different nodes in a quantum supercomputer could employ fundamentally different quantum technologies (such as superconducting circuits, trapped ions, or neutral atoms) for their local computations, while QNPUs provide a communication interface across the system. The clean separation of concerns not only simplifies current system design and implementation, but also provides a practical pathway toward realizing the heterogeneous quantum systems envisioned by the quantum computing community.